OpenStudy (anonymous):

The polynomial y = x2 - 2x + 1 has a repeated factor. (It may help to graph the equation.) True false

OpenStudy (anonymous):

I think it is false. i might be wrong i am not sure.

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

https://www.desmos.com/calculator/fo2tidmuyn If it has a repeated factor, then it crosses the x-axis at ONE point, does it?

OpenStudy (anonymous):

@oscar55773 What did you get?

OpenStudy (anonymous):

it crossed the x axis once

Directrix (directrix):

y = x2 - 2x + 1 = (x-1)^2 The repeated factor is of multiplicity two and indicates a point of tangency of the curve to the x-axis.